A small block slides without friction down an inclined starting form rest. Let Sn be the distance travelled from time


t = n-1 to t = n. Then Sn/Sn+1 is -


(A)   2n-1/2n


(B)   2n+1/2n+1


(C)   2n-1/2n+1


(D)   2n/2n+1

Grade:9

4 Answers

AKASH GOYAL AskiitiansExpert-IITD
13 years ago

Dear Shivam

S=ut+0.5at2= 0.5at2   (since u=0)

At t=n-1

S= 0.5a(n-1)2

At t=n

S'=0.5a(n)2

Sn= S'-S =0.5 [n2 - (n-1)2] = 0.5a (n-n+1)*(n+n-1)

Sn= 0.5a*(2n-1)

put n+1 in place of n in expression of Sn

Sn+1= 0.5a (2(n+1)-1)= 0.5a (2n+1)

Sn/Sn+1 = 2n-1/2n+1

option (C)

 

All the best.                                                           

AKASH GOYAL

AskiitiansExpert-IITD

 

Please feel free to post as many doubts on our discussion forum as you can. We are all IITians and here to help you in your IIT JEE preparation.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..

 

 


vikas askiitian expert
13 years ago

a is the accleration of block then

 S = 1/2 a t2                         (using this eq)

Sn = 1/2 a [n2-(n-1)2]

     = 1/2 a [2n-1]             ...........1

Sn+1 = 1/2a[(n+1)2-n2]

         =1/2 a [2n+1]          .........2

dividing eq 1 & 2

Sn/Sn+1 = 2n-1/2n+1

navkirat
6 years ago
Sir it is inclined plane so how do u took a=g as it is not horizontal plane So akash sir plz correct it
Kushagra Madhukar
3 years ago
Dear student,
Please find the solution to your problem.
 
S = ut+0.5at2 = 0.5at2   (since u=0)
At t = n – 1
S = 0.5a(n – 1)2
At t = n
S' = 0.5a(n)2
Sn = S' – S = 0.5 [n2 – (n – 1)2] = 0.5a (n – n + 1)*(n + n – 1)
Sn= 0.5a*(2n – 1)
put n+1 in place of n in expression of Sn
Sn+1 = 0.5a (2(n + 1) – 1) = 0.5a (2n+1)
Sn/Sn+1 = 2n – 1/2n + 1
Hence option (C) is correct
 
Thanks and regards,
Kushagra